OpenStudy (anonymous):

Give the general solution for : Tan (2x) + tan (x) = 1

OpenStudy (anonymous):

This is the answers from WolframAlpha C[1] Integers x= -2.83998+6.28319 C[1] x= -2.0026+6.28319 C[1] x= -0.976957+6.28319 C[1] x= 0.301616+6.28319 C[1] x= 1.13899+6.28319 C[1 x= 2.16464+6.28319 C[1])

OpenStudy (anonymous):

@siddarth95 Is it clear

OpenStudy (anonymous):

medal If I'm right

OpenStudy (anonymous):

I was hoping for the process to reduce it to the form of x=n(pi) + y (angle)

OpenStudy (anonymous):

you dont get medal for coppying answers !! i think so!! :)

OpenStudy (anonymous):

\[x=\tan^{-1} \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

oh no no ignore it!!

OpenStudy (anonymous):

i did 1 step wrong!!

OpenStudy (anonymous):

@deena123 The function you applied is truly wrong, @siddarth95 do you expect a elaborated answer

OpenStudy (anonymous):

\[\tan^3x-\tan^2x-3\tan x+1=0\]

OpenStudy (anonymous):

If not elaborate atleast to the level of know what to simplify

OpenStudy (anonymous):

Dude, question is a good one to try for IIT's here's the solution : open tan2x as first answer and convert it into an equation of tanx put tan x = y then take y to R.H.S to make it 1-y open the denominator of L.H.S. as 1-y^2 = (1+y) (1-y) take 1-y to R.H.S to make it (1-y)^2 then 1-y = +L.H.S. or -L.H.S. solve for both cases independently. Source: http://in.answers.yahoo.com/question/index?qid=20110318113330AACstbi

OpenStudy (anonymous):

@siddarth95 Its wat u wanted